OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -12, -7, -2, ...

OpenStudy (anonymous):

find a and b then calculate Tn

OpenStudy (anonymous):

@msumner the equation would be an = -17 + 5(n + 2) ?

OpenStudy (anonymous):

@rayanlovessarah how do I do that?

OpenStudy (anonymous):

@msumner thats not an answer choice

OpenStudy (anonymous):

an = -17 + 5(n + 2) an = -17 + 5(n + 1) an = -17 + 5(n - 1) an = -17 x 5(n - 1) these are the choices

OpenStudy (anonymous):

sorry -17 + d = -12 -17 + 12 = -d -5 = -d 5 = d \[a_n=a_1+(n-1)d\] where a1 is the first term of the sequence, d is the common difference, n is the number of the term to find. \[a_n=-17+(n-1)5\] which is the same as \[a_n = -17 + 5(n - 1)\]

OpenStudy (anonymous):

write down the formula to find any term of an arithmetic sequence

OpenStudy (anonymous):

@msumner thanks